What is the loan payment for a $128,000 home bought with a 20% down payment and the balance financed for 30 years at 9.5%
Answer by jim_thompson5910(35256)

20% down payment, so 80% is financed

80% of $128,000 = 0.8*128000 = 102,400 is financed (ie loaned out)

Turn to the formula P = L[c(1 + c)^n]/[(1 + c)^n - 1]

P = monthly payment
L = amount loaned out
c = monthly interest rate
n = number of payments

In this case

P = unknown
L = 102400
c = 0.095/12 = 0.00791666666666667
n = 30*12 = 360

Plug all that into the formula to get

P = L[c(1 + c)^n]/[(1 + c)^n - 1]

P = 102400[0.00791666666666667(1 + 0.00791666666666667)^360]/[(1 + 0.00791666666666667)^360 - 1]

P = 861.03470815104

P = 861.03

The monthly payment is $861.03